Suppose T€L(R^2) is defined by T(x, y) =(-3y,x).find the eigenvalues of T


1
Expert's answer
2021-06-29T12:07:52-0400


Let λ be an eigen value of T

T such that T(x,y) = λ(x,y)



Then we have T(x,y) = λ(x,y)


T(x,y) = λ(x,y) 

So we have


(−3y,x) = (λx,λy)


Thus 


λx = −3y

λx = −3y and x = λy

So we have 


λ(λy) = −3y

Which implies we have λ2 = - 3


which has no real solution. So there aren't any real eigen values but has imaginary eigen values.


"\lambda" 2 = -3,

"\lambda" = "i\sqrt{3}, \, -i\sqrt{3}" .